23. Map out the correct chromosome map for the following genes: (please draw out)

X-Z = 18% X-G = 35% Z-Y = 27% Z-G = 53% X-Y = 45%

A. Y-Z-X-G

B. X-Z-G-Y

C. Y-X-Z-G

D. Z-G-Y-X
Answer #1

Answer: Option 1 is correct
Explanation:
Recombination frequency is directly proportional to the distance between two genes.
Greater the distance between two genes, there is a greater chance of recombination between them.

Given information:
X-Z = 18% X-G = 35% Z-Y = 27% Z-G = 53% X-Y = 45%

From the given data, linkage map can be constructed as

Y--(27%)--Z--(18%)--X--(35%)--G